OpenStudy (anonymous):

Using the graph below, calculate the average rate of change for f(x) from x = 2 to x = 6. please help me solve

Directrix (directrix):

>>Using the graph below, Please post the graph.

OpenStudy (michele_laino):

hint: the requested average rate \(r\), is given by the subsequent formula: \[r = \frac{{f\left( 6 \right) - f\left( 2 \right)}}{{6 - 2}} = ...?\]